OpenStudy (anonymous):

Write the unit rate. A car traveled 330 miles in 6 hours. How fast was the car traveling? (mph) PLZ HELP lol idk unit rate :l

OpenStudy (alexandervonhumboldt2):

330/6

OpenStudy (danjs):

\[\frac{ 330~mi }{ 6~hr }=\frac{ x~mi }{ 1~hr }\]

OpenStudy (danjs):

m.p.h =\[\frac{ mi }{ }\]

OpenStudy (anonymous):

ty

OpenStudy (godlovesme):

@K3trina28 can u close the question if u r done with this post please? :) thanks!